OpenStudy (anonymous):

(1/27)^x = 9^x-4 Hey (: I'm lost can someone help me figure this out?

OpenStudy (freckles):

I bet you mean (1/27)^x=9^(x-4)

OpenStudy (freckles):

if so we can write both sides with base 3

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[27=3^3 \\ \text{ so } \frac{1}{27}=3^{-3}\]

OpenStudy (freckles):

and I think you write 9 as 3^to something

OpenStudy (freckles):

after you do that just set exponents equal and solve for x

OpenStudy (anonymous):

cool thank you!

OpenStudy (freckles):

you can show me any steps or show me what you get and I can check it

OpenStudy (anonymous):

I got 0=-4

OpenStudy (freckles):

\[(\frac{1}{27})^x=9^{x-4} \\ (3^{-3})^x=3^{2(x-4)} \\ 3^{-3x}=3^{2(x-4)}\] so you did this right?

OpenStudy (freckles):

and then since you had the same base on both sides you just needed to solve -3x=2(x-4) for x

OpenStudy (anonymous):

oh I tried x=x-4

OpenStudy (freckles):

no you have to get the same base on both sides before you can set exponents equal

OpenStudy (freckles):

\[a^x=a^y \text{ \implies } x=y \]

OpenStudy (anonymous):

8/5?

OpenStudy (freckles):

that sounds totally better

OpenStudy (anonymous):

haha thank you for your help! :)

OpenStudy (freckles):

np